OpenStudy (anonymous):

Juan took Brooke out for dinner at the Outback. The bill was $40. Juan left a 15% tip. Using 'T' as the total of the bill and the tip, write an algebraic equation that describes this transaction.

OpenStudy (johnweldon1993):

so the total of the bill would be 40 + whatever tip he left right? so he left a 15 % tip......what is 15% of 40?

OpenStudy (anonymous):

T=40(0.15)

OpenStudy (johnweldon1993):

well no...it would actually be T = 40 + 40(.15) if it was just T=40(.15) that would only be the tip that was left...you have to add in the meal price so T = 40 + 40(.15)

OpenStudy (anonymous):

but in real life don't you do 40(.15) I dont get where is the other +40 coming from?

OpenStudy (anonymous):

which the tip is $6

OpenStudy (johnweldon1993):

" The bill was $40. Juan left a 15% tip" the bill was 40.......that our original cost....... with a 15% tip...to find that...we do 40(.15) right? this gives us our tip... our total "T" would be our original cost...+ our tip so T = 40 + 40(.15)

OpenStudy (johnweldon1993):

and plus remember the last time you went to a restaurant....was the tip INCLUDED in the total of the bill? no, you get the bill, then ADD a tip :)

OpenStudy (anonymous):

oh yes.

OpenStudy (johnweldon1993):

understand it now? :)

OpenStudy (anonymous):

yes

OpenStudy (johnweldon1993):

great :) glad i could help!